# Causal Spatio-Temporal Prediction: An Effective and Efficient Multi-Modal Approach

### v1

@@ 2 Related Work
Multi-Modal Spatio-Temporal Prediction. In time series analysis, multi-modal methods have been applied across domains such as healthcare and finance. For instance, combining heterogeneous data sources (e.g., clinical records, medical images, genomic data) enhances medical predictive accuracy [58], while integrating social and economic signals strengthens market forecasting in finance [46]. Recent work explores LLM-based multi-modal models, GPT4MTS [20] leverages both numerical and textual inputs via prompt-based learning. MM-TSFlib [30] integrates language and time series models through end-to-end training. Wang et al. [48] combine LLMs with generative proxies for joint reasoning over news events and time series data, improving complex event prediction. However, these approaches focus solely on temporal dynamics and neglect spatial dependencies. In spatio-temporal prediction, Deng et al. [8] adopt self-supervised learning to capture latent patterns in multi-modal spatio-temporal data. Zhang et al. [62] apply AutoML techniques to model the dynamics of multi-modal meteorological data, but their method processes different traffic and weather types separately, with limited cross-modal interaction. Wang [44] utilizes multi-modal data for smart mobility prediction, yet each task relies on a single modality, falling short of unified multi-modal modeling. Zhao et al. [64] incorporate POI and weather data as contextual factors to model multi-modal traffic flow. Yan et al. [53] and Zhou et al. [67] fuse diverse urban data sources, including social media and real estate information, to improve traffic speed prediction. Han et al. [16] employ a pre-trained encoder and multi-modal inputs to model event impacts on traffic. Nonetheless, these models often struggle to model complex cross-modal dependencies and are mostly limited to textual inputs, lacking support for other essential data types like images.
@@ 3 Preliminary
Problem Statement (Multi-Modal Spatio-Temporal Prediction). Given the graph $\mathcal{G}$ and the historical $T$-step of multi-modal spatio-temporal data $\mathcal{X}$, where $\mathcal{X}$ may include spatio-temporal sequence and auxiliary modalities, we aim to learn a function $\theta(\cdot)$ that leverages $\mathcal{X}$ to predict the future $S$-step spatio-temporal data $Y_{\text{st}}=[y_{\text{st}}^{t+1},\dots,y_{\text{st}}^{t+S}]\in\mathbb{R}^{S\times N\times d}$.
@@ 4 Methodology
Framework Overview. As illustrated in Fig. 2, the E2-CSTP framework consists of three key components: (i) cross-modal feature fusion, (ii) dual-branch causal inference, and (iii) spatio-temporal encoding and decoding (STED). These modules work in concert to integrate multi-modal signals and apply causal inference techniques to reduce the impact of external confounders, thereby enhancing both prediction accuracy and model interpretability. Next, we detail each module below.
@@ 4.1 Cross-Modal Feature Fusion
[…]
@@ 4.2 Dual-Branch Causal Inference
To enhance model stability and robustness, we adopt an exponential moving average. The adjacency matrix is updated every $P$ epochs, effectively capturing dynamic changes in spatio-temporal data.
@@ 5.2 Overall Performance Comparison (RQ1)
First, E2-CSTP consistently outperforms all baselines across metrics and datasets, with MAE improvements ranging from 1.61% to 9.66% over the second-best methods. This highlights the effectiveness of our E2-CSTP framework in capturing fine-grained dependencies often missed by other methods.
Second, multi-modal and causality-based models generally outperform single-modal baselines, confirming the necessity of auxiliary modalities and causal reasoning in spatio-temporal forecasting. However, causality-based models, despite their theoretical appeal, often lack the contextual understanding required for comprehensive predictions. In contrast, E2-CSTP integrates causal reasoning within a multi-modal framework, further reducing uncertainty through cross-modal interactions.
@@ 5.4 Model Efficiency Study (RQ3)
To assess the effectiveness of our prediction module STED on computational cost, we replace it with several popular Transformer-based alternatives, including Informer [68], Autoformer [50], FEDformer [69], and iTransformer [31]. These variants are denoted as w/ In, w/ Auto, w/ FED, and w/ iTrans, respectively. Fig. 5 shows the prediction accuracy and per-epoch runtime on the Terra dataset. As observed, our method improves prediction accuracy by 1.78%–5.45% while reducing computational overhead by 17.37%–56.11% compared to these alternatives, demonstrating that the proposed module is both effective and efficient in practice.
@@ B.1 The definition of causal inference
Causal inference seeks to quantify how interventions influence outcomes of interest. A Structural Causal Model (SCM) specifies the underlying data-generating mechanism. The do-operator supplies the formal notation for interventions. Identifiability criteria, most notably the backdoor and frontdoor rules, convert causal queries into estimable statistical functionals. Pearl’s [37] three rules of do-calculus unify these components, allowing one to strip away interventions iteratively whenever the requisite d-separation conditions are met.
B.2 The three rules of do-calculus
Pearl’s do-calculus provides symbolic rules that transform interventional distributions to observational ones by exploiting graphical separation statements. Let $G_{do(x)}$ be the graph obtained from SCM after performing $do(X=x)$. $G_{\mathrm{null}(z)}$ denotes the graph where incoming edges to $Z$ are cut but $Z$ is not fixed to a constant.
@@ B.3 Deriving the backdoor adjustment
Consider a treatment–outcome pair $(X,Y)$ and a set of variables $Z$ such that (i) $Z$ blocks every backdoor path from $X$ to $Y$ in $\mathcal{G}$, and (ii) $Z$ contains no descendant of $X$. We recover the celebrated backdoor formula by successive applications of the do-calculus rules.
@@ B.4 Validity of the adjustment mechanism
Eq. 21 states that, whenever a valid back-door set $Z$ exists, causal effects can be estimated by (i) stratifying on $Z$, (ii) computing the conditional association between $X$ and $Y$ within each stratum, and (iii) averaging the results over the marginal distribution of $Z$. This principle underlies the dual-branch causal adjustment in Section 4.2 of the main paper, where the latent confounder $S$ plays the role of $Z$.
@@ Appendix C Training
2$\triangleright$ Preprocessing: Align modalities to obtain $\widetilde{X}_{\text{text}}$, $\widetilde{X}_{\text{img}}$ aligned with $X_{\text{st}}$;
@@ D.6 Parameter Sensitivity Study
Specifically, for Terra and GreenEarthNet, a lower $\lambda$ (0.25) works better, which places greater emphasis on the causal graph while reducing the weight of the raw spatial structure. This suggests that in these datasets, dominated by stable and consistent spatial patterns, relying more on causal refinement improves robustness and avoids redundancy. In contrast, BjTT and BikeNYC benefit from a more balanced integration (0.5), where the influence of the raw spatial and causal graphs is equally weighted. This indicates that both types of structural information are important for these datasets, likely due to the presence of more dynamic, heterogeneous, or noisy spatial-temporal patterns in urban environments.
@@
For Terra, the optimal $\beta$ is 0.75, highlighting the importance of spatio-temporal modeling in this dataset. GreenEarthNet and BjTT achieve optimal performance with a $\beta$ of 0.5, suggesting a balanced contribution of both branches. In contrast, BikeBYC benefits from a $\beta$ of 0.75. Although this dataset does not include multi-modal inputs, the auxiliary branch still encodes causally refined spatio-temporal features, which contribute to improved prediction performance.

### v2

@@ 2 Related Work
Multi-Modal Spatio-Temporal Prediction. In time series analysis, multi-modal methods have been applied across domains such as healthcare and finance. For instance, combining heterogeneous data sources (e.g., clinical records, medical images, genomic data) enhances medical predictive accuracy [59], while integrating social and economic signals strengthens market forecasting in finance [47]. Recent work explores LLM-based multi-modal models, GPT4MTS [21] leverages both numerical and textual inputs via prompt-based learning. MM-TSFlib [31] integrates language and time series models through end-to-end training. Wang et al. [49] combine LLMs with generative models for joint reasoning over news events and time series data, improving complex event prediction. However, these approaches focus solely on temporal dynamics and neglect spatial dependencies. In spatio-temporal prediction, Deng et al. [9] adopt self-supervised learning to capture latent patterns in multi-modal spatio-temporal data. Zhang et al. [63] apply AutoML techniques to model the dynamics of multi-modal meteorological data, but their method processes different traffic and weather types separately, with limited cross-modal interaction. Wang [45] utilizes multi-modal data for smart mobility prediction, yet each task relies on a single modality, falling short of unified multi-modal modeling. Zhao et al. [65] incorporate POI and weather data as contextual factors to model multi-modal traffic flow. Yan et al. [54] and Zhou et al. [68] fuse diverse urban data sources, including social media and real estate information, to improve traffic speed prediction. Han et al. [17] employ a pre-trained encoder and multi-modal inputs to model event impacts on traffic. Nonetheless, these models often struggle to model complex cross-modal dependencies and are mostly limited to textual inputs, lacking support for other essential data types like images.
@@ 3 Preliminary
Problem Statement (Multi-Modal Spatio-Temporal Prediction). Given the graph $\mathcal{G}$ and the historical $T$-step of multi-modal spatio-temporal data $\mathcal{X}$, where $\mathcal{X}$ may include spatio-temporal sequence and auxiliary modalities, we aim to learn a function $\theta(\cdot)$ that leverages $\mathcal{X}$ to predict the future $S$-step spatio-temporal data $Y_{\text{st}}=[y_{\text{st}}^{t+1},\dots,y_{\text{st}}^{t+S}]\in\mathbb{R}^{S\times N_{st}\times d}$.
@@ 4 Methodology
Framework Overview. As illustrated in Fig. 2, the E2‑CSTP framework consists of three key components: (i) cross-modal feature fusion, (ii) dual-branch causal inference, and (iii) spatio-temporal encoding and decoding (STED). These modules work in concert to integrate multi-modal signals and apply causal inference techniques to reduce the impact of external confounders, thereby enhancing both prediction accuracy and model interpretability. Next, we detail each module below.
@@ 4.1 Cross-Modal Feature Fusion
$\mathbf{M}^{t}_{i,j}=\begin{cases}1,&\text{if }j=\underset{k}{\arg\min}|\tau^{(i)}-\tau^{(k)}_{\text{st}}|\\ 0,&\text{otherwise}\end{cases},\mathbf{M}^{s}_{i,j}=\begin{cases}1,&\text{if }j=\underset{k}{\arg\min}\|\rho^{(i)}-\rho^{(k)}_{\text{st}}\|_{2}\\ 0,&\text{otherwise}\end{cases},$ (1)
@@ 4.2 Dual-Branch Causal Inference
To enhance model stability and robustness, we adopt an exponential moving average and update the adjacency matrix every $P=5$ epochs. This interval, determined empirically, balances stability and computational efficiency. Updating too frequently introduces noisy structural fluctuations, while a moderate update interval allows the model to refine its internal representations and capture meaningful spatio-temporal dependencies more effectively.
@@ 4.4 Model Training
In addition, the proposed STED module, whose complexity scales with sequence length $T$ and graph size $N$, drives the speed-up, while the shared text and image encoders merely handle preprocessing and therefore do not affect the relative ranking.
@@ 5.2 Overall Performance Comparison (RQ1)
First, E2‑CSTP consistently outperforms all baselines across metrics and datasets, with MAE improvements ranging from 1.61% to 9.66% over the second-best methods. This highlights the effectiveness of our E2‑CSTP framework in capturing fine-grained dependencies often missed by other methods.
Second, multi-modal and causality-based models generally outperform single-modal baselines, confirming the necessity of auxiliary modalities and causal reasoning in spatio-temporal forecasting. However, causality-based models, despite their theoretical appeal, often lack the contextual understanding required for comprehensive predictions. In contrast, E2‑CSTP integrates causal reasoning within a multi-modal framework, further reducing uncertainty through cross-modal interactions.
@@ 5.4 Model Efficiency Study (RQ3)
To assess the effectiveness of our prediction module STED on computational cost, we replace it with several popular Transformer-based alternatives, including Informer [69], Autoformer [51], FEDformer [70], and iTransformer [32], thus isolating the gain brought by replacing a quadratic Transformer with our linear GCN‑Mamba block. These variants are denoted as w/ In, w/ Auto, w/ FED, and w/ iTrans, respectively. Fig. 5 shows the prediction accuracy and per-epoch runtime on the Terra dataset. As observed, our method improves prediction accuracy by 1.78%–5.45% while reducing computational overhead by 17.37%–56.11% compared to these alternatives, demonstrating that the proposed module is both effective and efficient in practice.
@@ 6 Conclusion
7 Acknowledgment
This work was supported in part by the NSFC under Grants No. (62402422, 62025206, U23A20296, and 62472377), Yongjiang Talent Introduction Programme (2024A-162-G), Zhejiang Provincial Natural Science Foundation of China under Grant No. LZ25F020001, and Zhejiang Province’s "Lingyan" R&D Project under Grant No. 2024C01259. Ziquan Fang is the corresponding author.
@@ B.1 The definition of causal inference
Causal inference seeks to quantify how interventions influence outcomes of interest. A Structural Causal Model (SCM) specifies the underlying data‑generating mechanism. The do‑operator supplies the formal notation for interventions. Identifiability criteria, most notably the backdoor and frontdoor rules, convert causal queries into estimable statistical functionals. Pearl’s [38] three rules of do‑calculus unify these components, allowing one to strip away interventions iteratively whenever the requisite d‑separation conditions are met.
B.2 The three rules of do‑calculus
Pearl’s do‑calculus provides symbolic rules that transform interventional distributions to observational ones by exploiting graphical separation statements. Let $G_{do(x)}$ be the graph obtained from SCM after performing $do(X=x)$. $G_{\mathrm{null}(z)}$ denotes the graph where incoming edges to $Z$ are cut but $Z$ is not fixed to a constant.
@@ B.3 Deriving the backdoor adjustment
Consider a treatment–outcome pair $(X,Y)$ and a set of variables $Z$ such that (i) $Z$ blocks every backdoor path from $X$ to $Y$ in $\mathcal{G}$, and (ii) $Z$ contains no descendant of $X$. We recover the celebrated backdoor formula by successive applications of the do‑calculus rules.
@@ B.4 Validity of the adjustment mechanism
Eq. 21 states that, whenever a valid back‑door set $Z$ exists, causal effects can be estimated by (i) stratifying on $Z$, (ii) computing the conditional association between $X$ and $Y$ within each stratum, and (iii) averaging the results over the marginal distribution of $Z$. This principle underlies the dual‑branch causal adjustment in Section 4.2 of the main paper, where the latent confounder $S$ plays the role of $Z$.
@@ Appendix C Training
[…]
@@ D.6 Parameter Sensitivity Study
Specifically, for Terra and GreenEarthNet (remote sensing tasks), a lower $\lambda$ (0.25) works better, which places greater emphasis on the causal graph while reducing the weight of the raw spatial structure. These datasets contain stable spatial correlations (e.g., land use or vegetation types), making the data-driven refinement more valuable than default proximity-based connections. This suggests that in these datasets, dominated by stable and consistent spatial patterns, relying more on causal refinement improves robustness and avoids redundancy. In contrast, BjTT and BikeNYC (urban mobility tasks) benefit from a more balanced integration (0.5), where the influence of the raw spatial and causal graphs is equally weighted. These datasets involve highly dynamic and noisy spatio-temporal flows, requiring a blend of prior structure and adaptive refinement. This indicates that both types of structural information are important for these datasets, likely due to the presence of more dynamic, heterogeneous, or noisy spatial-temporal patterns in urban environments.
@@
For Terra, the optimal $\beta$ is 0.75, highlighting the importance of spatio-temporal modeling in this dataset. GreenEarthNet and BjTT achieve optimal performance with a $\beta$ of 0.5, suggesting a balanced contribution of both branches. In contrast, BikeBYC benefits from a $\beta$ of 0.75. Although this dataset does not include multi-modal inputs, the auxiliary branch still encodes causally refined spatio-temporal features, which contribute to improved prediction performance. Similar reasoning applies to $\beta$, where datasets with stronger exogenous influence (e.g., weather, events) benefit from greater auxiliary-branch emphasis, while those with purer internal spatio-temporal structure favor the primary prediction path.